The length of a rectangular parking lot is 10 feet longer than twice its width. The area of the lot is 3,600 sq. ft. Write and solve an equation for the length of the parking lot.
I wrote the equation (2W+10)*W = 3,600 sq. ft. but I do not know how to solve it.

Answer by moou_b(3)

ANSWER.
the equation is (2w+10)*w = 3600
2w^2+10w = 3600
w^2+5w = 1800 (dividing both sides by 2)
w^2+5w-1800 = 0
w = (-b +- sqrt( b^2-4*a*c ))/(2*a)
w = (-5 +- sqrt( 5^2-4*1*(-1800))/(2*1)
w = (-5 +- sqrt( 25+7200)/2
w = (-5 +- sqrt (7225))/2
w = (-5 +- 85)/2

taking the '+' sign, w1= (-5+85)/2
w1= 80/2 = 40
taking the '-' sign, w2= (-5-85)/2
w2= -90/2 = -45
therefore,length of the parking lot is 40 feet.